### v2

# Localizing Before Answering: A Hallucination Evaluation Benchmark for Grounded Medical Multimodal LLMs

Abstract
Medical Large Multi-modal Models (LMMs) have demonstrated remarkable capabilities in medical data interpretation. However, these models frequently generate hallucinations contradicting source evidence, particularly due to inadequate localization reasoning. This work reveals a critical limitation in current medical LMMs: instead of analyzing relevant pathological regions, they often rely on linguistic patterns or attend to irrelevant image areas when responding to disease-related queries. To address this, we introduce HEAL-MedVQA (Hallucination Evaluation via Localization MedVQA), a comprehensive benchmark designed to evaluate LMMs’ localization abilities and hallucination robustness. HEAL-MedVQA features (i) two innovative evaluation protocols to assess visual and textual shortcut learning, and (ii) a dataset of 67K VQA pairs, with doctor-annotated anatomical segmentation masks for pathological regions. To improve visual reasoning, we propose the Localize-before-Answer (LobA) framework, which trains LMMs to localize target regions of interest and self-prompt to emphasize segmented pathological areas, generating grounded and reliable answers. Experimental results demonstrate that our approach significantly outperforms state-of-the-art biomedical LMMs on the challenging HEAL-MedVQA benchmark, advancing robustness in medical VQA.
1 Introduction
Large Multimodal Models (LMMs), or Multi-modal Large Language Models (M-LLMs) Alayrac et al. (2022); Bai et al. (2023); Liu et al. (2024b) such as GPT-4V OpenAI et al. (2023) achieve superb performance in understanding data from multiple modalities, such as vision, and language, and generating human-like texts. Leveraging the strong capabilities of those foundational models, recent works Li et al. (2024); Chen et al. (2024); Zhang et al. (2023a); Chen et al. (2023); Moor et al. (2023); Wu et al. (2023b, a); Yan and Pei (2022) have developed M-LLMs for medical imaging applications such as medical Visual Question Answering (VQA). Recently, many multimodal foundation models (LLaVA-Med Li et al. (2024), CheXagent Chen et al. (2024), GPT-4 OpenAI et al. (2023), Gemini Team et al. (2023)) have emerged and demonstrated impressive reasoning and comprehension capabilities on biomedical queries and related domains. These models can assist clinical professionals in interpreting and gaining insights on many medical images, helping the diagnosis and treatment processes be more efficient.
Figure 1: An example of Med-VQA from our dataset. We follow LVLM-Interpret Stan et al. (2024) to visualize a) the attention map of the image with respect to the model response, and b) the relevancy score between the answer and input words. LLaVA-Med Li et al. (2024), as reference model fails to attend to the region of interest, and its answer has a low relevancy score to the image. Several hallucination reduction methods can generate responses more relevant to the image, but may not localize the target region.
Due to the sensitive nature of the biomedical domain, the accuracy and trustworthiness of these foundational models are important. Though impressive, Large Language Models (LLMs) and multimodal LLMs are prone to hallucinate contents. Specifically, the models produce false answers, which are not grounded in the visual evidence. Given a question such as “Does the patient have pneumonia in his right lung?”, many LLMs are prone to answer “Yes” without performing a deep analysis of the image content. This behavior arises because the observation pneumonia frequently co-occurs with parts of the lung, leading to shallow correlations rather than rigorous reasoning. Fig. 1 illustrates the answer of some open-source LMMs, as well as the visual cross-attention weight of the models during generation. These models tend to hallucinate and attend to irrelevant image regions when answering these questions.
This study shows two types of shortcut learning that current multimodal LLMs suffer from: textual and visual shortcut learning. First, the models attend to non-important text tokens with higher relevancy score than any image tokens. As shown in Fig. 1b, the model produces higher relevancy scores on token lung than any image token. Second, on the image, the model attends to non-queried image regions (as shown in Fig. 1a). This shows the model learns shortcut on irrelevant text and image tokens, instead of basing on the visual evidence when answering the questions.
To address these concerns, we introduce a new benchmark, dubbed HEAL-MedVQA (Hallucination EvAluation via Localization in Medical VQA) to dissect the LLM’s ability to localize at visual evidence when answering. Specifically, this paper proposes two evaluation protocols, called Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT), which probe LLMs’ sensitivity towards textual and visual shortcuts. To support this evaluation, we curate adversarial VQAs and procure pixel-level annotations of affected anatomy regions from three board-certified radiologists from two large-scale public datasets, MIMIC-CXR Johnson et al. (2019) and VinDr-CXR Nguyen et al. (2020). In total, we construct over 60,000 question-answer pairs probing LLM’s textual and visual shortcut learning. Our curated medical benchmark provides structured codebase, facilitating comprehensive comparisons of advanced LLMs.
To this end, we propose a Localize-before-Answer (LobA) framework that forces LMM to generate visual-evidence tokens, which is then input to segmentation decoder to localize the area of interests before answering. Based on the localization map, LobA self-prompts to enhance the model’s attention, forcing it to leverage visual evidence relevant to the question before answering. In summary, our main contributions are as follows.
  • We introduce HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA) benchmark consists of 67K QA pairs, and doctor-annotated segmentation masks with two new evaluation protocols, assessing LLMs’ robustness to textual and visual shortcut learning. Our benchmark includes a well-structured codebase, and comprehensive comparisons of 8 state-of-the-art LLMs and hallucination techniques, thus standardizing hallucination and accuracy evaluations of future multi-modal LLMs.
  • We propose the Localize-before-Answer (LobA) framework, which forces the model to localize the affected anatomical regions and self-prompt to enhance the attention on the target regions before answering.
  • Our proposed LobA framework significantly surpasses recent LLMs by up to 5.44%, while showing the highest robustness to textual and visual shortcut learning.
2 Related works
Medical Visual Question Answering. Several benchmarks in Med-VQA propose specialized datasets Hasan et al. (2018); Abacha et al. (2019, 2020, 2021). VQA-RAD Lau et al. (2018) offers both open-ended and closed-ended questions. SLAKE Liu et al. (2021) develops bilingual dataset with knowledge graph. Overall, existing benchmarks use accuracy as the main metrics, which overlooks the hallucination evaluation of multi-modal LLMs.
A majority of Med-VQA models are finetuned from other VLMs, namely LLaVA-Med Li et al. (2024) from LLaVA Liu et al. (2024b), CheXagent Chen et al. (2024) from BLIP-2 Li et al. (2023a), or Zhang et al. (2023a); Chen et al. (2023); Moor et al. (2023); Wu et al. (2023b, a); Yan and Pei (2022). Recent approaches, which localize relevant regions in Med-VQA have gained attention in recent times Tascon-Morales et al. (2023, 2024); Zhang et al. (2024a). However, during inference time, these VQA models require human-annotated masks to localize region of interests. In contrast, our method explicitly learns to localize during training, bypassing the needs to acquire human annotations when inferring.
Hallucination in VQA. To overcome hallucination, there are two main approaches. First, data-centric methods focus on augmenting existing VQA training datasets and propose new benchmarks for evaluation. In Li et al. (2023b) and Liu et al. (2023), they show that frequently appeared objects tend to guide models’ answers during inference time more than absent or less frequent ones. As such, POPE Li et al. (2023b) adds adversarial questions asking for the presence of objects that are absent on the image. LRV-Instruction Liu et al. (2023) introduces negative instructions which are irrelevant to the visual content.
Second, model-centric methods focus on modifying the model architecture to support the visual component. Intern-VL Chen and others (2024), and LLaVA-1.5-HD Liu et al. (2024a) scale up the vision foundation model to allow better visual understanding. Others, such as He et al. (2024); Jain et al. (2024) append several levels of information (segmentation mask, object list, task-based knowledge) to enrich the visual context. Lastly, several methods proposes inference-based methods, which augments the inference of the pre-trained LLMs. VCD Leng et al. (2024), and M3ID Favero et al. (2024) aim to ground inference by negating a contrastive version of the input where the image is distorted or dropped. OPERA Huang et al. (2024) discourages overly confident but erroneous information accumulated in self-attention layers by applying a penalty. Overall, existing hallucination techniques are applied on the image level, while neglecting region-level details. In contrast, we propose to explicitly reason on the pixel level before answering.
3 Hallucination Evaluation Benchmark
This section introduces HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA) benchmark consists of 67K QA pairs, and doctor-annotated segmentation masks with two new evaluation protocols.
3.1 New Evaluation Protocol — Sensitivity Test
Most current VQA works Abacha et al. (2019); Liu et al. (2021); Lau et al. (2018) primarily rely on the accuracy of the model’s output answers as the key evaluation metric. However, these metrics overlook other bias factors that could influence the model’s responses. In particular, we aim to assess two key phenomena in medical LMMs:
  • Language Shortcut: Despite generating correct answers, the model relies on language shortcut learnt in the training dataset, rather than the actual visual content. For example, given a question “Does the Right lower lung have lung opacity”, models can answer “Yes” due to the co-occurence of the words “lung” and “lung opacity” in the training dataset.
  • Vision Shortcut: The model might concentrate on non-relevant regions (as illustrated in Fig. 1a). This behavior results in the unreliability of LMMs, even when providing correct answers, as they fail to rely on medical visual evidence.
To that end, we propose two new evaluation protocols called Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT), as shown in Fig. 2.
Figure 2: Textual and Visual Perturbation Tests for our experiments. For Textual Perturbation Test (TPT), we randomly replace either the anatomy or the disease in the original question with another anatomy/disease so that the new disease pair is logical (present in the training dataset). For Visual Perturbation Test (VPT), we replace the visual area of interest (e.g. left lower lung) with the segmented area of the same region from another image that does not share the same diagnosis as the current one. We measure the percentage of instances a model changes the response from “Yes” to “No” to evaluate the model’s sensitivity to visual and language shortcuts.
Textual Perturbation Test. In disease diagnosis, medical LLMs can exploit the co-occurrence of disease and commonly affected regions to provide superficial answers. To assess how sensitive a model is to these language biases, we propose to swap the key entities, i.e., anatomy or disease, and evaluate the model’s sensitivity to the perturbed questions. The evaluation protocol is as follows.
  • From a test set of Yes/No binary questions under the form of “Does anatomy have disease”, we collect the True Positive samples of the model (samples correctly predicted as “Yes”).
  • We randomly replace either the anatomy or disease with another new anatomy/or disease. The new anatomy and disease pair has to be logical, which means they does co-occur in the dataset. For example, as shown in Fig. 2 (Above), the question “Does left lower lung suffer from pneumonia?” can be replaced by either switching the anatomy to right lower lung, or the disease to pulmonary fibrosis.
  • We measure the percentage of the answers being changed from “Yes” to “No” when applying this protocol as the textual perturbation test (TPT) score.
Visual Perturbation Test. We propose to assess whether the model grounds the answer on the anatomical region of interests. Specifically, when querying the presence of a disease in a particular area, e.g., the left lower lung, we take a segmented region of the left lower lung from another image that does not share the same diagnosis as the current one. This segmented region is resized and blended onto the original image. By doing so, we expect the model to focus its attention on the replaced region of interest to produce the correct diagnosis, altering the response from “Yes” to “No.” We define the visual perturbation test (VPT) score as the percentage of answers that change from “Yes” to “No” under this protocol.
3.2 Dataset Curation
To assess the LLM’s robustness against both textual and visual shortcuts, we introduce a new benchmark called HEAL-MedVQA (Hallucination Evaluation via Localization in Medical VQA). Built upon the two large scale radiology datasets, VinDr-CXR Nguyen et al. (2020) and MIMIC-CXR Johnson et al. (2019), we construct HEAL-MedVQA by mapping spatial relationships between diseases and anatomical locations, and probing the model’s robustness towards disease and location perturbation on both texts and images. Fig. 4 illustrates the overall pipeline of our data construction. We aim to extract the disease bounding boxes, and anatomical masks from the image. Based on the overlapping between disease and anatomy localization, we determine their spatial relationships. Then several QA templates are used to generate VQA samples.
Figure 3: Data processing pipeline. Our data processing pipeline consists of 4 main steps: 1. Anatomy Segmentation., where we obtain segmentation masks of core anatomies; 2. Disease Extraction., where the disease labels and their bounding boxes are obtained through either a Detection model, or from radiologists’ annotation; 3. Anatomy-Disease Relation Mapping, where anatomies are mapped with diseases based on their IoU score; 4. QA Generation, where question-answer pairs are generated based on the anatomy-disease mappings.
For the disease bounding boxes, VinDr-CXR database provides bounding box annotations for 22 medical findings. We then train a standard detection model, YOLOv5 Jocher et al. (2020), on VinDR-CXR dataset and obtain the bounding boxes on MIMIC-CXR. To reduce the false positive detection, we extract the disease labels from the associated medical reports, and filter out diseases which are absent on that patient.
For anatomy segmentation, we define eight main anatomies, e.g. Right lower lung, Left middle lung, Aorta, Heart, Trachea, where chest diseases commonly occur. On the training dataset, we then obtain the segmentation mask by using a model pre-trained on PAXRay++ dataset Seibold and others (2022). On the test datasets, we consult three board-certified radiologists to annotate 8 main anatomical structures on the pixel level. By providing doctor-annotated pixel-level segmentation, this research offers a new benchmark to evaluate the visual reasoning and the robustness to hallucination of current Multimodal LLMs.
To get the anatomy-disease relation, we measure the overlapping between the disease bounding box and anatomical segmentation mask. If the IoU score between a disease-anatomy pair is over a threshold $\delta=0.5$, then that disease occurs on the mentioned anatomy.
After obtaining a list of disease-anatomy mappings, for each image we randomly generate 2-5 questions and answers pairs based on these mappings. We focus on two main question formats: Close-ended questions and Open-ended questions, as exemplified by Fig. 4 Close-ended questions are binary questions of formats such as “Does anatomy have disease”, “Is there disease present in anatomy”. Based on the disease-anatomy mappings, the close-ended questions is either of type Positive, which queries the disease that is present in the anatomy, or of type Hallucinated, which queries the disease that is not present in the anatomy.
Open-ended questions require the model to makes more comprehensive diagnosis. It includes two types of questions: normal and abnormality-query questions which respectively probe regions without and with diseases. Sample question templates include “Are there any abnormalities in the heart?” or “List all abnormalities in the Aorta”.
Figure 4: Four question types in our benchmark.
In total, our HEAL-MedVQA dataset comprises 45,331 closed-ended questions and 22,573 open-ended ones. We compare our benchmark with those previously reported at Table 1. To the best of our knowledge, we are the first benchmark providing large-scale dataset with comprehensive hallucination metrics, and doctor-annotated masks to evaluate the hallucination of emerging LLMs.
Datasets # Images # QA pairs
Adversarial
questions
Closed-ended Open-ended
Visual
grounding
Annotation
VQA-Rad Lau et al. (2018) 315 3.5K
SLAKE Liu et al. (2021) 642 14K Mask/Bbox
VQA-Med Abacha et al. (2021) 5.5K 5.5K
PMC-VQA Zhang et al. (2023b) 149K 227K
OmniMedVQA Hu et al. (2024) 118k 128K
CARES Xia et al. (2024) 18K 41K
Halt-MedVQA Wu et al. (2024) 1736 2359
ProbMed Yan et al. (2024) 6.3K 57K
HEAL-MedVQA (Ours) 34K 67K Mask
Table 1: Comparison of Med-VQA datasets. indicates that visual annotations are for all anatomic regions in the image instead of the one in the question.
4 Proposed Method
Figure 5: The proposed LoBA frameworks consists of two phases. Grounded text generation: during training, our model learns to localize and segment a region of interest mentioned in the question. Self-prompting: during inferencing, the model attention is calibrated to attend to the segmented region. With the highlighted attention, the model refines and generates visually grounded answers that are more robust to hallucinations.
This section presents our proposed simple-yet-effective framework, Localize-before-Answer (LobA) to (i) localizes the queried regions, and then (ii) reweigh the attention to ground on correct visual evidence before answering. An overview of the framework is shown in Figure 5.
4.1 Grounded Text Generation
Given an image $x_{img}$ and a question about it $x_{text}$. We train LLMs to generate an answer $\hat{y}_{text}$, while grounding on the queried region, and producing a segmentation mask $\hat{M}$. The mask $\hat{M}$ corresponds to both $x_{text}$ and $\hat{y}_{text}$, explaining the response of the model.
We use a <SEG> token, denoting the token that will be used in later steps to obtain the segmentation of the relevant anatomies. Using this token, we instruct LLMs to localize on visual evidence before answering. As such, we format the response as “The location of {relevant areas} is at <SEG>”. For example, if the question is “Are there any abnormalities at the right lower lung?”, the output of the LMM will be “The location of the right lower lung is at <SEG>”. Note that, in our format, the model has to produce the visual evidence first, before giving the final answers.
The LLM is trained to generate the above response:
$\hat{y}_{text}=LLM(x_{text},x_{img}).$ (1)
We use the projected hidden state $h_{seg}$ from the <SEG> token as the query to a the segmentation model. This can be formulated as $\hat{M}=Seg(x_{img},h_{seg})$. Our LLM is trained via the text objectives. Similar to other LLMs Touvron et al. (2023); OpenAI et al. (2023), we formulate the text generation loss as the cross entropy loss between the generated text and the ground truth
$\mathcal{L}_{text}=CE(\hat{y}_{text},y_{text}).$ (2)
We combine BCE loss and DICE loss to train the segmentation module:
$\mathcal{L}_{seg}=\lambda_{BCE}BCE(\hat{M},M)+\lambda_{DICE}DICE(\hat{M},M).$ (3)
Finally, the overall objective can be seen as the weighted sum of the two losses
$\mathcal{L}=\lambda_{text}\mathcal{L}_{text}+\lambda_{seg}\mathcal{L}_{seg}.$ (4)
4.2 Attention Highlighter via Self-Prompting
While the proposed training paradigm allows the model to localize the region of interests when answering, hallucination can still occur when the model pays little attention to the image compared to the text. To deal with this, we propose to highlight the attention on the queried regions by self-prompting the segmentation map. Our module consists of two functions: attention reweighting and contrastive decoding.
For attention reweighting, patches of interest $x_{hl}$ are added extra weight before feeding through the softmax layer so that those regions get more focus on during answer generation. Denote $h_{j,i}$ the original value of the attention logits of patch j to patch $i$, and $\beta$ the hyperparameter of added weight, then the new attention weight after reweighting $\hat{h}_{j,i}$ becomes
$\hat{h}_{j,i}=\begin{cases}h_{j,i}&\text{if }i\notin x_{hl},\\ h_{j,i}+\log\beta,\beta>0&\text{if }i\in x_{hl}.\end{cases}$ (5)
Then the attention probability after Softmax is
$\tilde{a}_{j,i}=\frac{\beta^{m_{i}}\exp{(a_{j,i})}}{\sum_{k}{\beta^{m_{k}}\exp{a_{j,k}}}}.$ (6)
Here, $m_{i}=1$, indicating token $i$ is highlighted, and $m_{i}=0$ otherwise. We apply this function to the visual backbone of M-LLM where the patches of interests are interpolated from the mask. Applying the reweighted attention $\tilde{A}$, we obtain the highlighted image tokens $\tilde{x}_{\text{img}}$:
$\tilde{x}_{\text{img}}=\tilde{A}V.$ (7)
To further alleviate shortcut learning, we perform contrastive decoding Zhang et al. (2024b), which contrasts the model decisions after and before highlighting regions of interests. Particularly, we compute the difference between the token probability after highlighting $p_{hl}$ and the token probability before highlighting $p_{bh}$ to obtain the decoded probability $p$:
$\begin{split}p_{\text{bh}}&=p(\hat{y}_{\text{text}}|x_{\text{text}},x_{\text{img}})\\ p_{\text{hl}}&=p(\hat{y}_{\text{text}}|x_{\text{text}},\tilde{x}_{\text{img}})\\ p&=\text{softmax}((1+\alpha)\log p_{\text{hl}}-\alpha\log p_{\text{bh}}).\end{split}$ (8)
where $\alpha$ is a hyperparameter controlling the degree of grounding. Applying contrastive decoding minimizes the effect of other tokens to the model output, thus preventing hallucination.
Model MIMIC VinDr
Yes/No Open-ended Yes/No Open-ended
F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall
Proprietary Models
Gemini-1.5-Flash-8B Team et al. (2023) 0.548 0.520 0.569 0.091 0.127 0.024 0.571 0.604 0.543 0.038 0.098 0.023
GPT-4-O OpenAI et al. (2023) 0.646 0.706 0.556 0.109 0.143 0.089 0.599 0.660 0.487 0.097 0.087 0.069
GPT-4-Vision OpenAI et al. (2023) 0.521 0.652 0.409 0.087 0.093 0.067 0.568 0.637 0.424 0.074 0.053 0.072
Open-source Models
CheXagent Chen et al. (2024) 0.677 0.621 0.695 0.425 0.450 0.416 0.610 0.628 0.603 0.374 0.342 0.330
BioMedGPT Zhang et al. (2023a) 0.708 0.734 0.672 0.397 0.387 0.412 0.624 0.606 0.631 0.289 0.265 0.310
MedFlamingo Alayrac et al. (2022) 0.562 0.574 0.552 0.365 0.390 0.310 0.584 0.600 0.547 0.301 0.285 0.322
LLaVA-Med Li et al. (2024) 0.722 0.712 0.709 0.551 0.541 0.565 0.703 0.708 0.699 0.523 0.519 0.530
LISA Lai et al. (2024) 0.716 0.692 0.728 0.524 0.556 0.511 0.707 0.712 0.694 0.521 0.509 0.534
Adversarial Techniques
PH + LLaVA-Med Zhang et al. (2024b) 0.735 0.754 0.724 0.563 0.589 0.540 0.716 0.725 0.698 0.522 0.510 0.527
VCD + LLaVA-Med Leng et al. (2024) 0.716 0.708 0.721 0.566 0.583 0.538 0.652 0.699 0.591 0.501 0.547 0.497
CRG + LLaVA-Med Zhang et al. (2022) 0.731 0.715 0.725 0.544 0.563 0.554 0.689 0.667 0.702 0.509 0.524 0.494
LobA w/o self-prompt 0.727 0.716 0.731 0.544 0.557 0.535 0.724 0.745 0.706 0.511 0.526 0.507
LobA (Ours) 0.752 0.759 0.743 0.581 0.593 0.572 0.728 0.758 0.711 0.542 0.533 0.560
Table 2: Performance comparison across different models on the proposed HEAL-MedVQA benchmark on MIMIC and VinDr datasets for Yes/No and Open-ended VQA tasks. Text in bold and underline highlights the best and second best results, respectively.
Discussion. Although Prompt Highlighter Zhang et al. (2024b) also emphasizes the “highlighting” of regions of interest to reduce hallucination, ours differ in two folds. First, Prompt Highlighter requires human prompting for attention highlighting, which is expensive in medical image analysis. In contrast, ours automatically localizes regions of interests, thus obviating the need for medical professionals to manually localize regions of interest. Second, Prompt Highlighter only enhances attention on the language decoder, our methodology reweighs attention scores on the visual backbone. The deep intervention on visual backbone is more effective in mitigating the visual shortcut learning.
5 Experimental Results
In this section, we report the benchmark results of popular medical LMMs and our very own framework, as well as further analysis for our curated adversarial grounding VQA benchmark.
MIMIC VinDr
Method Anatomy Disease Anatomy Disease
Gemini-1.5 0.532 0.510 0.521 0.487
GPT-4-O 0.594 0.603 0.509 0.596
CheXagent 0.621 0.650 0.689 0.650
BioMedGPT 0.687 0.610 0.623 0.612
LLaVA-Med 0.764 0.762 0.751 0.734
PH + LLaVA-Med 0.783 0.774 0.724 0.762
VCD + LLaVA-Med 0.744 0.788 0.740 0.750
CRG + LLaVA-Med 0.731 0.771 0.742 0.738
LobA (Ours) 0.792 0.801 0.770 0.748
Table 3: Textual Perturbation Test on MIMIC and VinDr datasets
MIMIC VinDr
Gemini-1.5 0.410 0.403
GPT-4-O 0.557 0.581
CheXagent 0.631 0.671
BioMedGPT 0.647 0.653
LLaVA-Med 0.696 0.680
PH + LLaVA-Med 0.720 0.698
VCD + LLaVA-Med 0.678 0.654
CRG + LLaVA-Med 0.682 0.693
LobA (Ours) 0.734 0.701
Table 4: Visual Perturbation Test on MIMIC and VinDr datasets
Module Backbone LobA LobA
w/o self-prompt (Ours)
Component Localization
Self-prompt
Perb Test TPT 0.763 0.772 0.797
VPT 0.743 0.738 0.759
F1 Score MIMIC (Yes/No) 0.722 0.727 0.752
MIMIC (Open-ended) 0.551 0.544 0.581
VinDr (Yes/No) 0.703 0.724 0.728
VinDr (Open-ended) 0.523 0.511 0.542
Table 5: Ablation studies of different components of LobA. The TPT and VPT scores are averages for MIMIC and VinDr datasets.
Benchmarks. We conduct a thorough and systemic evaluation of the most popular Medical Large Multimodal Models on our large-scale HEAL-MedVQA benchmark. Specifically, we utilize 7 state-of-the-art multi-modal LLMs for our evaluation protocol, 3 of which are proprietary models: GPT-4o, GPT-4 Vision OpenAI et al. (2023) and Gemini 1.5 Team et al. (2023) and 4 are open source M-LLMs of the medical domain: CheXagent Chen et al. (2024), LLaVA-Med Li et al. (2024), BioMedGPT Zhang et al. (2023a) and Med-Flamingo Alayrac et al. (2022). We also assess the capabilities of state-of-the-art adversarial VQA methods, including VCD Leng et al. (2024), Prompt Highlighter Zhang et al. (2024b), and CRG Wan et al. (2025).
Implementation Details. Open source multi-modal LLMs are fine-tuned on our training dataset with LoRA Hu et al. (2022) before evaluation. The learning rate is tuned from the range of values $\{1e-4,1e-5,1e-6\}$. For LobA framework, we used LLaVA-Med Li et al. (2024) as the pretrained MLLM and MedSAM Kirillov et al. (2023) as the segmentation model. Following Zhang et al. (2024b), we select $\alpha=0.3$, $\beta_{ViT}=\beta_{LLM}=2$ as the set of hyperparameters for the Attention Highlighter module. All models are fine-tuned using Transformer, PyTorch and DeepSpeed frameworks on a single 80GB A100 GPU cluster. During inference, the temperature of all multi-modal LLMs is set to 0.1, and the beam size is set to 1.
Evaluation Metrics. To assess the accuracy of the multi-modal LLMs, we utilize a powerful LLM, LLaMa 3.1-8B specifically, capable of extracting information from the M-LLM’s output, and categorizing them into a set of disease labels. For example, for an open-ended question, if the model returns the answer “The heart suffers from pneumonia, pulmonary fibrosis and nodule/mass” then the extracted labels by the LLM will be pneumonia, pulmonary fibrosis, nodule/mass. We report the multi-label precision, recall, and micro F1 score as the main accuracy metrics. For hallucination sensitivity analysis, we adopt our proposed evaluation metrics, Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT) scores, as discussed in Sec. 3.1.
[…]
Textual Perturbation Test. Table 3 reports the sensitivity of each model to textual language bias, measuring the percentage of changed True Positive answers when either switching an anatomy or a disease. The higher the percentage is, the more robust a model is from textual shortcut bias. Our proposed framework achieves the highest sensitivity measure overall, with the ratio of changed answers when changing anatomy/disease to be around 0.77 to 0.80. Enhancing the attention to the queried regions as in our proposal helps alleviate the textual and visual shortcut learning. Lastly, fine-tuned open-source models are generally more robust than closed-source models like GPT-4-O, showing that fine-tuning LLMs on our adversarial data remedies their hallucination problems.
[…]
Ablation studies. Table 5 reports the effects of each component in our LobA framework: the LLaVA Med backbone, Grounded Text Generation module, and the full framework with self-prompting. Applying Grounded Text Generation improves the model’s performance in some categories, showing the effectiveness of localization-aware training.With the self-prompting framework, the accuracy and textual/visual robustness improves greatly, most evidently from 0.763 to 0.797 in TPT score, or 0.722 to 0.752 in MIMIC-CXR dataset Yes/No F1 Score.
Qualitative analysis. Fig 11 showcases our LobA and some other LMM’s response on a few qualitative examples, as well as the visual attention map on the input image. We can see that LobA manages to query the attention weights much more intuitively compared to LLaVA-Med or LISA While LLaVA-Med and LISA’s attention are scattered into many irrelevant areas, in the first example LobA’s main attention weights are allocated around the right lower lung area, and the second example is near the left upper lung. Thanks to the grounding generation module, followed by LobA’s self prompting module to automatically re-weight the attention to the area of interest, the model is able to have better visual reasoning capabilities on the area of interest, leading to more accurate answers compared with other methods.
6 Conclusion
Figure 6: Qualitative case study of both question types
This paper introduces Heal-MedVQA a new large scale Medical Visual Question Answering Benchmark with over 67,000 question-answer pairs, which queries diseases presence at local anatomies, evaluating LMM’s capabilities to localize at grounded visual evidence when answering. In addition, we present the Localize-before-Answer (LobA) framework which trains the LMMs to segment the region of interest and re-adjust their attention for more emphasis on the segmented pathological areas, leading to more reliable answers. Our experimental results showed our framework LobA outperforming state-of-the-art medical LMMs on Heal-MedVQA, proving its robustness and localization capabilities.
Appendix A Computational Efficiency
We present the inference time (seconds) and GPU consumption (MB) of LoBA on a system with a single A100 80GB GPU for 100 samples in comparison with some other medical LMMs in Tab. 6. Our LobA yields computational efficiency comparable to that of existing methods, while achieving superior localization capabilities.
CheXagent Med-Flamingo LLaVA-Med LobA (Ours)
Inf. Time 113.4 126.5 87.0 101.2
GPU Mem 23406 26453 17729 25254
Table 6: Comparison of inference time and GPU consumption
Appendix B Benchmarking Other General LLMs
We report the F1 score of additional proprietary and open-source LMMs (fine-tuned) on the VinDr subset in Tab. 7
Model Gemini-1.5 Claude 3.7 Qwen 2.5 VL InternVL LobA
Pro Sonnet 7B-Instruct 2.5 8B (Ours)
Yes/No 0.583 0.610 0.660 0.654 0.728
Open 0.078 0.062 0.413 0.378 0.542
Table 7: Performance on VinDr for general LMMs
Appendix C Ablation study
We conducted two ablation studies, examining the effect of added attention weights $\beta$ (Tab. 8) and the effect of segmentation quality on the model’s results (Fig 7).
In summary, low weights cause the segmented area to have minimal influence, harming results, while excessive weight (e.g., 4.0) shifts focus away from the question, also degrading performance. For the second ablation study, the quality of the segmentation indeed correlates directly with LoBA’s performance. Overall, our model outperforms LISA baselines when the segmentation Dice score $>0.25$, which accounts for more than $10\%$ of the test set. This shows that our design is already beneficial even when the segmentation module can localize the rough regions to attend to. For samples with Dice $<0.25$, which only account for 9.47% of data, our LoBA yields a slight performance decrease.
$\beta$ Yes/No F1 Open F1
0.5 0.694 0.476
1 0.698 0.510
2 0.720 0.545
3 0.728 0.542
4 0.712 0.501
Table 8: Effect of $\beta$ on performance on VinDr
Figure 7: Yes/No F1 w.r.t segmentation qualities. Data portion (%) on test set is also shown.
Appendix D LobA Prompt Template
Figure 8: LobA Instruction Template
Figure 9: GPT and Gemini Models Instruction Template for Yes/No questions
Figure 10: GPT and Gemini Models Instruction Template for Open-ended questions
Figure 11: Qualitative Examples from our dataset
We propose a Grounded Text Generation training paradigm to enable LMM to localize the area of interest with the segmentation token “[SEG]“, before giving its answer for the model to be aware of the localization information.
As such, our model is fine-tuned to follow a strict template as “The location of {relevant areas} is at <SEG>. The answer is …”. To make sure the model’s output comply with this standard, we designed a few-shot instruction template for LobA so that the model can follow the desired template more accurately, displayed in Fig 8
Appendix E Proprietary Models Template
For GPT-4-O and GPT-4-Vision and Gemini-1.5-Flash-8B models, we modify the template of these proprietary models for yes/no questions and open-ended questions separately to allow them to follow our instructions correctly. These 2 templates are displayed by Fig 10 and 9
Appendix F Evaluation Template
Figure 12: Yes/No evaluation prompt template
Figure 13: Open Ended evaluation prompt template
We utilize LLaMA-3 as the LLM backbone to extract the labels from the model’s natural language answers with the following prompt templates, Fig 12 and Fig 13
Appendix G Question example and qualitative results
We present some qualitative examples, as well as some model’s response and their visual attention map in Fig 11.